RC meeting and Run Plan, Thursday Oct. 11

From HallCWiki
Jump to: navigation, search


  • Beam quality First 20.5 hours Last 3.5 hours
    • Bottom line: terrible first ~19 hours, situation improved markedly in the last ~5 hours
    • Current fluctuating between 40uA and 50uA
  • Progress
    • Added 23% of the minimum statistics on LH2 for the current run
    • Completed dummy for the current kinematic point

Requiring attention

  • HMS Dipole LN2 fill leve very high (essentially full)
    • Not urgent, small ice ball forming in non-critical location (we're watching it)
    • Cause: seems like JT-4 is stuck again
    • Will coordinate with Andy Kenyon early next week for opportunistic escorted access
    • Alarm disabled for now, re-enable when fixed. Instructions
  • Rotation to small angle: Complications when going to 6.50 deg tonight:
    • Risk tripping the rotation alarm: Holly on standby
    • Need new limits on ion chamber diffusors if we run above >=50uA (Jay Benesh confirmed this is the case), more info here.
    • SHMS rates expected to be close to detector limit (tracking efficiency), might need to limit beam current.
      • Would take care of above issue. Current estimate from Garth puts us at <55uA.

Other requests (non-urgent)

  • RadCon wants to put a small detector in the SHMS hut during the next opportunity
    • Needs few hours, communicate with Brad when next opportunity arrises
  • Hall D needs opportunistic 1shift of HallC down for beam studies, not pressing right now (needs to be done before end-of-run)
    • Unlikely this will happen in the next week

Plans - Physics

  • Continue until ~midnight with Q2=3. W=3.14 x=0.25 theta_Shms=9.42, hopefully will reach 50-60% of statistics goal
  • Change SHMS angle to 6.50 deg (close to software minimum) for next kinematic point.
    • Aim for 55% of minimum statistics (same as 12 degree point)
    • Expect to stay there for ~5-7 shifts.